Q1: In the circuit shown below use Thevenins method to find Voc and the maximum power delivered to the load RL 23 Τον Θ
Answer #1

- (Queit) In the circuit shown below use theveninn method to find Voc and the maximum power delivered to the load Red 3r Lovapplying rodal analysin at node Ű Voc - 10 + Voc 48=0 Voc - 10+ 2V0C + 32 = 0 8 Voc + 22 = 0 voce -24 = -7-33 volt Vth = VOCparallel so resistance 4r 4 zu equivalent resistance are connected in Rx = 4x2 = 1.33 4+2 Now, resistance Rx f or are connect